# Guiding Cross-Modal Representations with MLLM Priors via Preference Alignment

### v1

@@ 1 Introduction
Cross-modal retrieval, which aims to retrieve relevant content across different modalities (e.g., retrieving images with text queries), has long been a core topic in the vision-language research community. It also serves as a fundamental building block for a wide range of downstream task, including visual question answering, retrieval-augmented generation, and increasingly, LLM-based multi-modal agent systems. Despite the remarkable success of large-scale contrastive pretraining frameworks such as CLIP [1], a fundamental challenge still remains: the modality gap, i.e., the discrepancy in feature representation between visual and textual modalities. This modality gap often manifests as a spatial separation between image and text embeddings in the shared latent space, and also limits the effectiveness of learned representations in various retrieval tasks [2, 3, 4]. Prior works [2, 4] have shown that this misalignment might be attributed from architectural choices, training dynamics, and even input information imbalance.
While prior works mainly focused on measuring the modality gap using explicit embeddings [2, 3], the alignment properties of models that operate directly via logits, such as off-the-shelf Multimodal Large Language Models (MLLMs), remain less explored. To better understand the modality gap across different model architectures, we propose a unified metric based on the 1-Wasserstein Distance (WD) [5] that enables direct comparison between logit-based and embedding-based models. For MLLMs, we compute pairwise alignment scores using output logits (as detailed in Section 2), while for embedding models, we use cosine similarities. In both cases, we apply WD to measure the discrepancy between the similarities distributions. Intriguingly, through this quantitative comparison, we discover that MLLMs like Qwen2-VL [6, 7] exhibit strong inherent modality alignment capabilities even without relying on explicit embedding representations. Our analysis, illustrated in Appendix Figure 3, reveals that MLLMs demonstrate an emergent ability to align image and text inputs more effectively than CLIP.
@@
Recent efforts have fine-tuned MLLMs for retrieval by enabling cross-modal representation learning [8, 9, 10]. However, transitioning from generative architectures to MLLM-based retrievers often diminishes the inherent alignment capabilities of MLLMs, even after further fine-tuning. Motivated by the above observation, we aim to adapt MLLM for retrieval tasks while preserving their strong cross-modal alignment strength.
To this end, we propose MAPLE (Modality-Aligned Preference Learning for Embeddings), a novel framework that bridges the alignment capabilities of off-the-shelf MLLM with MLLM-based retrieval model. MAPLE transfers MLLM’s alignment capabilities to embedding spaces through two key dimensions: data-level preference construction and training-strategy-level preference alignment. At the data level, we retrieve top-K hard samples for each anchor and leverage MLLM to score their matching degree with the anchor’s corresponding cross-modal content. These alignment scores are then used to construct both pairwise preferences and listwise preferences for training. At the training level, we derive the Relative Preference Alignment (RPA) loss from Direct Preference Optimization (DPO) [11], specifically adapted for embedding models to achieve fine-grained cross-modal alignment. Through optimization with the RPA loss on these constructed preferences, as illustrated in Figure 1 c, our method effectively captures subtle distinctions between retrieved images under the same query.
@@
  • We validate our framework through comprehensive experiments on different benchmarks including general retrieval (e.g., COCO [12], Flickr30K [13]), fine-grained retrieval (e.g. Winoground [14], NaturalBench [15], MMVP [16], BiVLC [17]). The experimental results demonstrate the superiority and effectiveness of our approach in both general and fine-grained retrieval tasks.
@@ 2 Preliminaries and Notation
Given a batch of $N$ image-text pairs $\{(x_{i}^{\text{img}},x_{i}^{\text{txt}})\}_{i=1}^{N}$, unlike CLIP, which utilizes separate encoders to extract embeddings for each modality, we leverage the MLLM, a unified architecture, to obtain embeddings for each modality. Following the prior works [8, 10], we construct prompts using predefined templates such as “<text> Describe this text in one word:” for text and “<image> Describe this image in one word:” for image. These prompts are used to process the image-text pairs, then we extract the corresponding text embeddings $\{z_{i}^{\text{txt}}\}_{i=1}^{N}$ and image embeddings $\{z_{i}^{\text{img}}\}_{i=1}^{N}$.
@@
where $\tau$ is a temperature hyperparameter. Prior works [18, 19] established contrastive learning inherently employs a coarse-grained alignment strategy that uniformly pushes away all negative samples in the embedding space, with limited consideration of fine-grained semantic similarity between these samples. This uniform treatment struggles to establish nuanced discriminative boundaries, particularly for semantically similar negatives.
@@
The modality gap represents the misalignment between visual and textual feature distributions. Prior work [2] measured this using the average distance between mean embeddings: $\left\|\vec{\Delta}_{\text{gap}}\right\|=\left\|\mu_{\text{txt}}-\mu_{\text{img}}\right\|$, where $\mu_{\text{mod}}=\frac{1}{N}\sum_{i=1}^{N}z_{i}^{\text{mod}}$.
@@
However, this mean-based comparison overlooks the full distributional characteristics of similarities. Furthermore, it requires explicit embeddings, making it less suitable to measure the modality gap for the logits-based model. To capture distributional discrepancy more effectively, we employ the 1-Wasserstein Distance (WD) [5] to compare similarity distributions. For two distributions $\mathbb{P}_{A}$ and $\mathbb{P}_{B}$, WD is defined as:
@@
We measure the modality gap using WD on the Winoground-style [14] fine-grained dataset, where each test instance contains two images ($x_{0}^{\text{img}},x_{1}^{\text{img}}$) and two captions ($x_{0}^{\text{txt}},x_{1}^{\text{txt}}$), forming two matching pairs $(x_{0}^{\text{img}},x_{0}^{\text{txt}})$ and $(x_{1}^{\text{img}},x_{1}^{\text{txt}})$. We denote the sets of all $x_{0}^{\text{txt}}$, $x_{1}^{\text{txt}}$, $x_{0}^{\text{img}}$, and $x_{1}^{\text{img}}$ as $T_{0}$, $T_{1}$, $I_{0}$, and $I_{1}$ respectively. We consider it from two perspectives: the distributional gap $W(\mathbb{P}_{T_{0}I_{0}},\mathbb{P}_{T_{0}T_{0}})$, which quantifies the alignment between intra-modal $\mathbb{P}_{T_{0}T_{0}}$ and cross-modal $\mathbb{P}_{T_{0}I_{0}}$ similarity distributions (where a lower value indicates better alignment but risks representation collapse [20] when approaching zero), and the discriminative gap $W(\mathbb{P}_{T_{0}I_{0}},\mathbb{P}_{T_{0}I_{1}})$, which assesses the model’s ability to distinguish between matching and non-matching pairs (where a higher value indicates better discriminative ability). Figure 3 in the Appendix illustrates the modality gap comparison between CLIP and Qwen2-VL based on this unified metric.
@@ 3 Method
We initialize our model from a pretrained MLLM backbone to inherit its multi-modal alignment capability. Refer to the work [21], to convert the autoregressive MLLM into a discriminative retrieval paradigm, we make two small modifications: replacing the causal attention mask with bidirectional attention and adding mean-pooling over final hidden states to aggregate sufficient features for retrieval.
@@ 3.1 Preference Data Construction
The pipeline begins with an offline stage to prepare candidate sets. We first extract DINOv2 [22] embeddings from the image dataset and apply Semantic Deduplication (SemDeup) [23]—a clustering-based method—to filter near-duplicate samples. For each deduplicated image $x_{i}^{\text{img}}$ in a batch, we retrieve its top-$K$ nearest neighbors $\{\hat{x}_{j}^{\text{img}}\}_{j=1}^{K}$ from the gallery based on cosine similarity, forming an image candidate set $\mathcal{C}_{i}^{\text{img}}=\{x_{i}^{\text{img}}\}\cup\{\hat{x}_{j}^{\text{img}}\}_{j=1}^{K}$. To enrich caption diversity and create challenging negatives, we leverage the multi-image reasoning capability of an off-the-shelf MLLM. We prompt the MLLM with the image set $\mathcal{C}_{i}^{\text{img}}$ to generate discriminative captions that explicitly highlight inter-image differences. This process yields a corresponding text candidate set $\mathcal{C}_{i}^{\text{txt}}=\{x_{i}^{\text{txt}}\}\cup\{\hat{x}_{j}^{\text{txt}}\}_{j=1}^{K}$. Candidate generation details are provided in Appendix B.1.
@@
  • Listwise Preferences: Instead of breaking the ranking into independent pairs, we leverage the structure of the entire ranked list $(x_{i,r_{0}},x_{i,r_{1}},\dots,x_{i,r_{K}})$. This approach considers preferences within all possible suffixes of the list. Specifically, for each starting rank $k$ (from 00 to $K-1$), the item $x_{i,r_{k}}$ is treated as the preferred item relative to the set of all subsequent items $\{x_{i,r_{j}}\}_{j=k+1}^{K}$ in the suffix $(x_{i,r_{k}},\dots,x_{i,r_{K}})$. This captures the relative ordering across the whole list more directly than pairwise comparisons.
@@ 3.2 Preference Alignment
In contrast to the contrastive loss’s coarse-grained alignment approach, our goal is to establish nuanced discriminative boundaries leveraging fine-grained preference data. Drawing inspiration from Direct Preference Optimization (DPO) [11], which effectively fine-tunes LLMs to align with human preferences, we similarly aim to fine-tune our MLLM-based retrieval model to align with the sophisticated preference signals from off-the-shelf MLLMs.
@@
The memory-inefficiency issue can be alleviated by adopting a uniform prior $U$ for the reference model $\pi_{w}$, similar to CPO [24]. In this case, the reference model terms $\pi_{w}\left(y_{w}\mid x\right)$ and $\pi_{w}\left(y_{l}\mid x\right)$ cancel out in the log-ratio difference (up to a constant), eliminating the need for costly computations and storage associated with $\pi_{w}$. The simplified objective becomes:
@@
Alternatively, the Listwise RPA approach directly optimizes the model’s ability to align with the MLLM’s ranking using the listwise preference data. Inspired by PRO [25], this loss encourages the model to assign the highest similarity score to the top-ranked item within each suffix of the MLLM’s ranked list. For a text anchor $x_{i}^{\text{txt}}$ and its ranked image candidates $(x_{i,r_{0}}^{\text{img}},\dots,x_{i,r_{K}}^{\text{img}})$, the loss iterates through each possible top element $x_{i,r_{k}}^{\text{img}}$ (for $k$ from 00 to $K-1$) and maximizes the log-probability of this element being ranked highest among the suffix $\{x_{i,r_{j}}^{\text{img}}\}_{j=k}^{K}$, using a softmax over the model’s similarity scores $s_{ij}^{\text{txt2img}}$. To incorporate the fine-grained preference strength from the MLLM, each term in the sum (corresponding to a specific suffix starting at $k$) is weighted by the average preference margin assigned by the MLLM to $x_{i,r_{k}}^{\text{img}}$ over the subsequent items in that suffix:
@@ 4.1 Experimental Setup
We train MAPLE on a curated subset of the OpenImage dataset [26] (details in Appendix B.1). We evaluate its performance on standard general (MS-COCO [12], Flickr30K [13]) and fine-grained (Winoground [14], NaturalBench [15], MMVP [16], BiVLC [17]) retrieval benchmarks using standard metrics (e.g., Image/Text Recall@1, Image/Text scores). About the more fine-grained evaluation details, please refer to Appendix B.3. MAPLE is compared against strong CLIP-based [1, 27, 28, 29, 30] and MLLM-based [8, 10] retrieval models. We use LoRA for fine-tuning MAPLE MLLM-based embedding models. Comprehensive implementation details are provided in Appendix B.2.
@@ 4.2 Main Results
Table 1: Performance Comparison on General and Fine-grained Retrieval Tasks. Best results are in bold, second best are underlined. VladVA haven’t released the model weights, so “--” represents unavailable results.
@@
CLIP (ViT-L[1] 58.1 37.0 87.2 67.3 27.5 12.3 41.8 45.0
OpenCLIP (ViT-G/14[27] 66.3 48.8 91.5 77.8 32.0 12.8 46.2 46.5
SigLIP (so/14[28] 70.2 52.0 93.5 80.5 37.5 16.3 62.7 63.9
SigLIPv2 (g/16-2B[30] 72.8 56.1 95.4 86.0 39.8 17.0 65.5 68.7
EVA-CLIP (8B[29] 70.1 52.0 94.5 80.3 36.5 14.8 58.5 59.3
EVA-CLIP (18B[29] 72.8 55.6 95.3 83.3 35.8 15.0 58.7 61.2
@@
E5-V(LLaVA-Next-8B[8] 62.0 52.0 88.2 79.5 32.3 14.8 60.3 67.6
VladVA(Qwen2-VL-2B[10] 71.9 52.5 93.7 80.4 - - - -
VladVA(LLaVA-1.5-7B[10] 72.9 59.0 94.3 83.3 40.5 17.5 - -

### v3

@@ 1 Introduction
Cross-modal retrieval, which aims to retrieve relevant content across different modalities (e.g., retrieving images with text queries), has long been a core topic in the vision-language research community. It also serves as a fundamental building block for a wide range of downstream task, including visual question answering, retrieval-augmented generation, and increasingly, LLM-based multi-modal agent systems. Despite the remarkable success of large-scale contrastive pretraining frameworks such as CLIP [radford2021learning], a fundamental challenge still remains: the modality gap, i.e., the discrepancy in feature representation between visual and textual modalities. This modality gap often manifests as a spatial separation between image and text embeddings in the shared latent space, and also limits the effectiveness of learned representations in various retrieval tasks [liang2022mind, jiang2023understanding, schrodi2024two]. Prior works [liang2022mind, schrodi2024two] have shown that this misalignment might be attributed from architectural choices, training dynamics, and even input information imbalance.
While prior works mainly focused on measuring the modality gap using explicit embeddings [liang2022mind, jiang2023understanding], the alignment properties of models that operate directly via logits, such as off-the-shelf Multimodal Large Language Models (MLLMs), remain less explored. To better understand the modality gap across different model architectures, we propose a unified metric based on the 1-Wasserstein Distance (WD) [panaretos2019statistical] that enables direct comparison between logit-based and embedding-based models. For MLLMs, we compute pairwise alignment scores using output logits (as detailed in Section 2), while for embedding models, we use cosine similarities. In both cases, we apply WD to measure the discrepancy between the similarities distributions. Intriguingly, through this quantitative comparison, we discover that MLLMs like Qwen2-VL [wang2024qwen2, bai2025qwen2] exhibit strong inherent modality alignment capabilities even without relying on explicit embedding representations. Our analysis, illustrated in Appendix Figure 3, reveals that MLLMs demonstrate an emergent ability to align image and text inputs more effectively than CLIP.
@@
Recent efforts have fine-tuned MLLMs for retrieval by enabling cross-modal representation learning [jiang2024e5, lin2024mm, ouali2024discriminative]. However, transitioning from generative architectures to MLLM-based retrievers often diminishes the inherent alignment capabilities of MLLMs, even after further fine-tuning. Motivated by the above observation, we aim to adapt MLLM for retrieval tasks while preserving their strong cross-modal alignment strength.
To this end, we propose MAPLE (Modality-Aligned Preference Learning for Embeddings), a novel framework that bridges the alignment capabilities of off-the-shelf MLLM with MLLM-based retrieval model. MAPLE transfers MLLM’s alignment capabilities to embedding spaces through two key dimensions: data-level preference construction and training-strategy-level preference alignment. At the data level, we retrieve top-K hard samples for each anchor and leverage MLLM to score their matching degree with the anchor’s corresponding cross-modal content. These alignment scores are then used to construct both pairwise preferences and listwise preferences for training. At the training level, we derive the Relative Preference Alignment (RPA) loss from Direct Preference Optimization (DPO) [rafailov2023direct], specifically adapted for embedding models to achieve fine-grained cross-modal alignment. Through optimization with the RPA loss on these constructed preferences, as illustrated in Figure 1 c, our method effectively captures subtle distinctions between retrieved images under the same query.
@@
  • We validate our framework through comprehensive experiments on different benchmarks including general retrieval (e.g., COCO [lin2014microsoft], Flickr30K [young2014image]), fine-grained retrieval (e.g. Winoground [thrush2022winoground], NaturalBench [li2024naturalbench], MMVP [tong2024eyes], BiVLC [miranda2024bivlc]). The experimental results demonstrate the superiority and effectiveness of our approach in both general and fine-grained retrieval tasks.
@@ 2 Preliminaries and Notation
Given a batch of $N$ image-text pairs $\{(x_{i}^{\text{img}},x_{i}^{\text{txt}})\}_{i=1}^{N}$, unlike CLIP, which utilizes separate encoders to extract embeddings for each modality, we leverage the MLLM, a unified architecture, to obtain embeddings for each modality. Following the prior works [jiang2024e5, ouali2024discriminative], we construct prompts using predefined templates such as “<text> Describe this text in one word:” for text and “<image> Describe this image in one word:” for image. These prompts are used to process the image-text pairs, then we extract the corresponding text embeddings $\{z_{i}^{\text{txt}}\}_{i=1}^{N}$ and image embeddings $\{z_{i}^{\text{img}}\}_{i=1}^{N}$.
@@
where $\tau$ is a temperature hyperparameter. Prior works [jiang2023vision, wang2023boosting] established contrastive learning inherently employs a coarse-grained alignment strategy that uniformly pushes away all negative samples in the embedding space, with limited consideration of fine-grained semantic similarity between these samples. This uniform treatment struggles to establish nuanced discriminative boundaries, particularly for semantically similar negatives.
@@
The modality gap represents the misalignment between visual and textual feature distributions. Prior work [liang2022mind] measured this using the average distance between mean embeddings: $\left\|\vec{\Delta}_{\text{gap}}\right\|=\left\|\mu_{\text{txt}}-\mu_{\text{img}}\right\|$, where $\mu_{\text{mod}}=\frac{1}{N}\sum_{i=1}^{N}z_{i}^{\text{mod}}$.
@@
However, this mean-based comparison overlooks the full distributional characteristics of similarities. Furthermore, it requires explicit embeddings, making it less suitable to measure the modality gap for the logits-based model. To capture distributional discrepancy more effectively, we employ the 1-Wasserstein Distance (WD) [panaretos2019statistical] to compare similarity distributions. For two distributions $\mathbb{P}_{A}$ and $\mathbb{P}_{B}$, WD is defined as:
@@
We measure the modality gap using WD on the Winoground-style [thrush2022winoground] fine-grained dataset, where each test instance contains two images ($x_{0}^{\text{img}},x_{1}^{\text{img}}$) and two captions ($x_{0}^{\text{txt}},x_{1}^{\text{txt}}$), forming two matching pairs $(x_{0}^{\text{img}},x_{0}^{\text{txt}})$ and $(x_{1}^{\text{img}},x_{1}^{\text{txt}})$. We denote the sets of all $x_{0}^{\text{txt}}$, $x_{1}^{\text{txt}}$, $x_{0}^{\text{img}}$, and $x_{1}^{\text{img}}$ as $T_{0}$, $T_{1}$, $I_{0}$, and $I_{1}$ respectively. We consider it from two perspectives: the distributional gap $W(\mathbb{P}_{T_{0}I_{0}},\mathbb{P}_{T_{0}T_{0}})$, which quantifies the alignment between intra-modal $\mathbb{P}_{T_{0}T_{0}}$ and cross-modal $\mathbb{P}_{T_{0}I_{0}}$ similarity distributions (where a lower value indicates better alignment but risks representation collapse [jing2021understanding] when approaching zero), and the discriminative gap $W(\mathbb{P}_{T_{0}I_{0}},\mathbb{P}_{T_{0}I_{1}})$, which assesses the model’s ability to distinguish between matching and non-matching pairs (where a higher value indicates better discriminative ability). Figure 3 in the Appendix illustrates the modality gap comparison between CLIP and Qwen2-VL based on this unified metric.
@@ 3 Method
We initialize our model from a pretrained MLLM backbone to inherit its multi-modal alignment capability. Refer to the work [lee2024nv], to convert the autoregressive MLLM into a discriminative retrieval paradigm, we make two small modifications: replacing the causal attention mask with bidirectional attention and adding mean-pooling over final hidden states to aggregate sufficient features for retrieval.
@@ 3.1 Preference Data Construction
The pipeline begins with an offline stage to prepare candidate sets. We first extract DINOv2 [oquab2023dinov2] embeddings from the image dataset and apply Semantic Deduplication (SemDeup) [abbas2023semdedup]—a clustering-based method—to filter near-duplicate samples. For each deduplicated image $x_{i}^{\text{img}}$ in a batch, we retrieve its top-$K$ nearest neighbors $\{\hat{x}_{j}^{\text{img}}\}_{j=1}^{K}$ from the gallery based on cosine similarity, forming an image candidate set $\mathcal{C}_{i}^{\text{img}}=\{x_{i}^{\text{img}}\}\cup\{\hat{x}_{j}^{\text{img}}\}_{j=1}^{K}$. To enrich caption diversity and create challenging negatives, we leverage the multi-image reasoning capability of an off-the-shelf MLLM. We prompt the MLLM with the image set $\mathcal{C}_{i}^{\text{img}}$ to generate discriminative captions that explicitly highlight inter-image differences. This process yields a corresponding text candidate set $\mathcal{C}_{i}^{\text{txt}}=\{x_{i}^{\text{txt}}\}\cup\{\hat{x}_{j}^{\text{txt}}\}_{j=1}^{K}$. Candidate generation details are provided in Appendix B.1.
@@
  • Listwise Preferences: Instead of breaking the ranking into independent pairs, we leverage the structure of the entire ranked list $(x_{i,r_{0}},x_{i,r_{1}},\dots,x_{i,r_{K}})$. This approach considers preferences within all possible suffixes of the list. Specifically, for each starting rank $k$ (from 0 to K1K-1), the item $x_{i,r_{k}}$ is treated as the preferred item relative to the set of all subsequent items $\{x_{i,r_{j}}\}_{j=k+1}^{K}$ in the suffix $(x_{i,r_{k}},\dots,x_{i,r_{K}})$. This captures the relative ordering across the whole list more directly than pairwise comparisons.
@@ 3.2 Preference Alignment
In contrast to the contrastive loss’s coarse-grained alignment approach, our goal is to establish nuanced discriminative boundaries leveraging fine-grained preference data. Drawing inspiration from Direct Preference Optimization (DPO) [rafailov2023direct], which effectively fine-tunes LLMs to align with human preferences, we similarly aim to fine-tune our MLLM-based retrieval model to align with the sophisticated preference signals from off-the-shelf MLLMs.
@@
The memory-inefficiency issue can be alleviated by adopting a uniform prior $U$ for the reference model $\pi_{w}$, similar to CPO [xu2024contrastive]. In this case, the reference model terms $\pi_{w}\left(y_{w}\mid x\right)$ and $\pi_{w}\left(y_{l}\mid x\right)$ cancel out in the log-ratio difference (up to a constant), eliminating the need for costly computations and storage associated with $\pi_{w}$. The simplified objective becomes:
@@
Alternatively, the Listwise RPA approach directly optimizes the model’s ability to align with the MLLM’s ranking using the listwise preference data. Inspired by PRO [song2024preference], this loss encourages the model to assign the highest similarity score to the top-ranked item within each suffix of the MLLM’s ranked list. For a text anchor $x_{i}^{\text{txt}}$ and its ranked image candidates $(x_{i,r_{0}}^{\text{img}},\dots,x_{i,r_{K}}^{\text{img}})$, the loss iterates through each possible top element $x_{i,r_{k}}^{\text{img}}$ (for $k$ from 0 to K1K-1) and maximizes the log-probability of this element being ranked highest among the suffix $\{x_{i,r_{j}}^{\text{img}}\}_{j=k}^{K}$, using a softmax over the model’s similarity scores $s_{ij}^{\text{txt2img}}$. To incorporate the fine-grained preference strength from the MLLM, each term in the sum (corresponding to a specific suffix starting at $k$) is weighted by the average preference margin assigned by the MLLM to $x_{i,r_{k}}^{\text{img}}$ over the subsequent items in that suffix:
@@ 4.1 Experimental Setup
We train MAPLE on a curated subset of the OpenImage dataset [kuznetsova2020open] (details in Appendix B.1). We evaluate its performance on standard general (MS-COCO [lin2014microsoft], Flickr30K [young2014image]) and fine-grained (Winoground [thrush2022winoground], NaturalBench [li2024naturalbench], MMVP [tong2024eyes], BiVLC [miranda2024bivlc]) retrieval benchmarks using standard metrics (e.g., Image/Text Recall@1, Image/Text scores). About the more fine-grained evaluation details, please refer to Appendix B.3. MAPLE is compared against strong CLIP-based [radford2021learning, schuhmann2022laion, zhai2023sigmoid, sun2024eva, tschannen2025siglip] and MLLM-based [jiang2024e5, ouali2024discriminative] retrieval models. We use LoRA for fine-tuning MAPLE MLLM-based embedding models. Comprehensive implementation details are provided in Appendix B.2.
@@ 4.2 Main Results
Table 1: Performance Comparison on General and Fine-grained Retrieval Tasks. Best results are in bold, second best are underlined. VladVA haven’t released the model weights, so “-” represents unavailable results.
@@
CLIP (ViT-L[radford2021learning] 58.1 37.0 87.2 67.3 27.5 12.3 41.8 45.0
OpenCLIP (ViT-G/14[schuhmann2022laion] 66.3 48.8 91.5 77.8 32.0 12.8 46.2 46.5
SigLIP (so/14[zhai2023sigmoid] 70.2 52.0 93.5 80.5 37.5 16.3 62.7 63.9
SigLIPv2 (g/16-2B[tschannen2025siglip] 72.8 56.1 95.4 86.0 39.8 17.0 65.5 68.7
EVA-CLIP (8B[sun2024eva] 70.1 52.0 94.5 80.3 36.5 14.8 58.5 59.3
EVA-CLIP (18B[sun2024eva] 72.8 55.6 95.3 83.3 35.8 15.0 58.7 61.2
@@
E5-V(LLaVA-Next-8B[jiang2024e5] 62.0 52.0 88.2 79.5 32.3 14.8 60.3 67.6
VladVA(Qwen2-VL-2B[ouali2024discriminative] 71.9 52.5 93.7 80.4 - - - -
VladVA(LLaVA-1.5-7B[ouali2024discriminative] 72.9 59.0 94.3 83.3 40.5 17.5 - -